Solve the differential equation by substitution suggested by equation. Show complete solution.


(x + 2y - 1)dx + 3(x + 2y)dy = 0


1
Expert's answer
2021-10-06T12:16:46-0400

Solution:

"(x + 2y - 1)dx + 3(x + 2y)dy = 0" ...(i)

Put "x+2y=v"

"\Rightarrow dx+2dy=dv"

Using these, (i) becomes,

"(v-1)(dv-2dy)+3vdy=0
\\\Rightarrow vdv-2vdy-dv+2dy+3vdy=0
\\\Rightarrow (v-1)dv+(v+2)dy=0
\\\Rightarrow (1-v)dv=(v+2)dy
\\\Rightarrow \dfrac{1-v}{v+2}dv=dy"

"\\\Rightarrow \dfrac{1}{v+2}dv-\dfrac{v}{v+2}dv=dy
\\\Rightarrow \dfrac{1}{v+2}dv-\dfrac{v+2-2}{v+2}dv=dy
\\\Rightarrow \dfrac{1}{v+2}dv-dv+\dfrac{2}{v+2}dv=dy
\\\Rightarrow \dfrac{3}{v+2}dv-dv=dy"

On integrating both sides,

"3 \ln|v+2|-v+c=y
\\\Rightarrow 3 \ln|x+2y+2|+x+2y+c=y
\\\Rightarrow 3 \ln|x+2y+2|+x+y+c=0"

which is required solution.